# To Write or to Automate Linguistic Prompts, That Is the Question

### 26 Mar 2026 v1

@@ Abstract
LLM performance is highly sensitive to prompt design, yet whether automatic prompt optimization can replace expert prompt engineering in linguistic tasks remains unexplored. We present the first systematic comparison of hand-crafted zero-shot expert prompts, base DSPy signatures, and GEPA-optimized DSPy signatures across translation, terminology insertion, and language quality assessment, evaluating five model configurations. Results are task-dependent. In terminology insertion, optimized and manual prompts produce mostly statistically indistinguishable quality. In translation, each approach wins on different models. In LQA, expert prompts achieve stronger error detection while optimization improves characterization. Across all tasks, GEPA elevates minimal DSPy signatures, and the majority of expert–optimized comparisons show no statistically significant difference. We note that the comparison is asymmetric: GEPA optimization searches programmatically over gold-standard splits, whereas expert prompts require in principle no labeled data, relying instead on domain expertise and iterative refinement.
11footnotetext: * Equal contribution.
@@ 1 Introduction
Large language models (LLMs) have shown remarkable capabilities across a wide range of natural language processing (NLP) tasks. However, research shows that LLM performance is highly sensitive to the way prompts are worded and structured. Sclar et al. [28] found that open-source LLMs exhibit significant performance differences when tested under slight, meaning-preserving variations in prompt formatting, such as adding or omitting punctuation marks; Voronov et al. [33] found that variations in how examples are selected, verbalized or separated in few-shot settings results in performance differences, while Mizrahi et al. [21] found that both manual and automated prompt paraphrasing lead to drastic performance variations in both open-source and closed models. While metrics have been proposed to evaluate this fragility [3, 28], improving prompts in response to such diagnoses remains a largely manual, iterative process requiring task-specific knowledge. This practice, referred to as prompt engineering, is the subject of a substantial body of research, producing for example catalogs of prompt patterns [34] and prompting techniques [27].
@@
We investigate this question across three tasks: (a) terminology insertion, where an existing translation is corrected to conform to a provided glossary; (b) translation; and (c) language quality assessment (LQA), where translation errors are detected and characterized using an MQM-compliant [19] schema.
@@ 2 Related Work
The well-documented fragility of LLM performance under prompt variation [28, 33, 21] has motivated a growing body of work on automatic prompt optimization. Early approaches treated the prompt as a learnable vector in the model’s embedding space, optimizing it via gradient-based [17, 18] or reinforcement learning [39] techniques. These methods require access to model gradients, ruling out black-box API models, and produce continuous representations that are not interpretable as natural language. Discrete approaches operate directly over natural language tokens, producing human-readable prompts. Earlier discrete methods still relied on gradient access [31, 30, 2, 38, 25], but recent work has shown that effective discrete optimization is possible without access to model parameters [36, 7, 1], making automatic optimization directly applicable to the black-box API models on which language service providers (LSPs) typically rely.
@@
We evaluate the manual-versus-automatic comparison on three linguistic tasks—machine translation (MT), language quality assessment (LQA), and terminology enforcement—where LSPs increasingly rely on LLMs and where prompt design directly affects production quality.
LLMs have progressed rapidly on MT: early evaluations showed competitive performance on high-resource European languages but gaps in low-resource settings [11, 9], with GPT-4 surpassing supervised baselines such as NLLB [32] in many directions while still lagging behind commercial systems in low-resource languages [41]. More recent benchmarks show LLMs now dominate, with Gemini 2.5 Pro placing in the top evaluation cluster for 14 of 16 language pairs [14]. Fine-grained human evaluations suggest LLM translations remain comparable to junior and mid-level professional translators but fall short of senior professionals [35].
@@ 3.1 Shared Experimental Setup
For all tasks, we use gold-standard annotations by expert linguists. Data is split into train (20%), validation (40%), and test (40%) sets, as in Agrawal et al. [1], who allocate the majority of data to validation and test since GEPA’s reflective optimization requires relatively few training examples. Test set sizes are 1085 examples for terminology insertion, 989 for translation, and 1250 for LQA. Per-language results are omitted for space reasons but can be found in the accompanying GitHub repository.
@@ 3.3 Translation
We optimize for the average between HTER and chrF3, as defined by Popovic et al [23]. We use the SacreBLEU implementation for both scores. HTER is a word-level distance metric while chrF3 is a character-level n-gram metric. Together, these metrics provide a strong comparison between the machine-translated text and human-translated reference. To normalize both scores to the same scale, we use the formula $score=\frac{100-HTER}{200}+\frac{chrF3}{100}$
@@ 5 Conclusions
GEPA optimization consistently elevates minimal DSPy signatures, often closing the gap to expert-level performance, but the degree and direction of remaining differences are task-dependent. In terminology insertion, the two approaches are largely statistically indistinguishable on translation quality (BLEU, HTER), with optimization producing significantly higher term match rates for several models but not consistently improving fluency. In translation, most model–prompt differences are not statistically significant, with the exception of Gemini 3.1 Flash-Lite, where expert prompts significantly outperform optimized ones on BLEU. In LQA, optimization yields statistically significant gains over manual prompts on error characterization for several models, more consistently for Gemini 3.1 Flash-Lite. The degree to which optimization improves over base DSPy signatures also varies by task: gains are largest and most consistent in translation and LQA, while terminology insertion shows smaller improvements.
@@
Unified single-stage systems generally match or outperform decomposed multi-stage pipelines—most clearly for error detection in LQA (0.05–0.20 points on Detection F1) and edit distance in terminology insertion—suggesting that practitioners should prefer end-to-end designs unless there is a specific reason to decompose. Separately, Qwen3:8B-q4_K_M, while performing worse than proprietary models on translation and terminology insertion, narrowed the gap on several LQA metrics, with optimization over base DSPy producing the largest single gain observed across all tasks and models.
@@
This study has several limitations. First, the comparison involves an inherent asymmetry in how each approach uses labeled data. GEPA optimization programmatically exploits gold-standard training and validation splits through automated feedback loops, whereas expert prompt engineering draws on domain knowledge, iterative testing, and selective inspection of examples, if available. Comparable performance between the two approaches does not indicate equivalent cost, as each demands a different balance of compute, labeled data, and human effort. Furthermore, the two approaches are not mutually exclusive: a domain expert could use GEPA optimization as a starting point and manually refine the resulting prompts, or conversely, use expert-crafted prompts as seed inputs for automatic optimization. Our study treats them as independent conditions, but a combined workflow may outperform either in isolation. Second, several design choices were scoped for practical reasons: we evaluate a single optimizer (GEPA) in its “light” mode with the dspy.Predict module, we do not evaluate few-shot or chain-of-thought manual prompts, and all evaluation is automatic without human judgments. Each of these choices may underestimate the ceiling of either approach — stronger optimizers, richer expert baselines, and human evaluation could all shift the results — but they were necessary to keep the study tractable across three tasks, five models, and multiple conditions.

### 05 May 2026 v2

@@ Abstract
LLM performance is highly sensitive to prompt design, yet whether automatic prompt optimization can replace expert prompt engineering in linguistic tasks remains unexplored. We present the first systematic comparison of hand-crafted zero-shot expert prompts, base DSPy signatures, and GEPA-optimized DSPy signatures across terminology insertion, translation and language quality assessment, evaluating five model configurations. Results are task-dependent. For terminology insertion and translation, GEPA-optimized prompts are competitive with expert prompts: most differences are not statistically significant, and optimization significantly improves glossary term match rates for several models. In language quality assessment, expert prompts achieve stronger error detection while optimization improves characterization. Across tasks, GEPA elevates minimal DSPy signatures, often closing the gap to expert performance. We note that the comparison is asymmetric: GEPA optimization searches programmatically over gold-standard splits, whereas expert prompts require in principle no labeled data, relying instead on domain expertise and iterative refinement.
11footnotetext: Equal contribution.
@@ 1 Introduction
LLMs have shown remarkable capabilities across a wide range of NLP tasks. However, research shows that LLM performance is highly sensitive to the way prompts are worded and structured. Sclar et al. [29] found that open-source LLMs exhibit significant performance differences when tested under slight, meaning-preserving variations in prompt formatting, such as adding or omitting punctuation marks; Voronov et al. [33] found that variations in how examples are selected, verbalized or separated in few-shot settings results in performance differences, while Mizrahi et al. [21] found that both manual and automated prompt paraphrasing lead to drastic performance variations in both open-source and closed models. While metrics have been proposed to evaluate this fragility [3, 29], improving prompts in response to such diagnoses remains a largely manual, iterative process requiring task-specific knowledge. This practice, referred to as prompt engineering, is the subject of a substantial body of research, producing for example catalogs of prompt patterns [34] and prompting techniques [28].
@@
We investigate this question across three linguistic tasks where language service providers (LSPs) increasingly rely on LLMs and where prompt design directly affects production quality: (a) terminology insertion, where an existing translation is corrected to conform to a provided glossary; (b) translation; and (c) language quality assessment (LQA), where translation errors are detected and characterized using an MQM-compliant [19] schema.
@@ 2 Related Work
The well-documented fragility of LLM performance under prompt variation [29, 33, 21] has motivated a growing body of work on automatic prompt optimization. Early approaches treated the prompt as a learnable vector in the model’s embedding space, optimizing it via gradient-based [17, 18] or reinforcement learning [39] techniques. These methods require access to model gradients, ruling out black-box API models, and produce continuous representations that are not interpretable as natural language. Discrete approaches operate directly over natural language tokens, producing human-readable prompts. Earlier discrete methods still relied on gradient access [32, 31, 2, 38, 26], but recent work has shown that effective discrete optimization is possible without access to model parameters [36, 7, 1], making automatic optimization directly applicable to the black-box API models on which LSPs typically rely.
@@
LLMs have progressed rapidly on machine translation: early evaluations showed competitive performance on high-resource European languages but gaps in low-resource settings [11, 9], with GPT-4 surpassing supervised baselines such as NLLB [23] in many directions while still lagging behind commercial systems in low-resource languages [41]. More recent benchmarks show LLMs now dominate, with Gemini 2.5 Pro placing in the top evaluation cluster for 14 of 16 language pairs [14]. Fine-grained human evaluations suggest LLM translations remain comparable to junior and mid-level professional translators but fall short of senior professionals [35].
@@ 3.1 Shared Experimental Setup
For all tasks, we use proprietary datasets spanning various domains, with gold-standard annotations produced by professional contract linguists. Annotations are reviewed internally, with in-house feedback used to refine and correct them before inclusion in the dataset. Data is split into train (20%), validation (40%), and test (40%) sets, as in Agrawal et al. [1], who allocate the majority of data to validation and test since GEPA’s reflective optimization requires relatively few training examples. Test set sizes are 1085 examples for terminology insertion, 989 for translation, and 1250 for LQA. Per-language results are omitted for space reasons but can be found in the accompanying GitHub repository.
@@ 3.3 Translation
We optimize for the average between HTER and chrF3, as defined by Popović et al [24]. We use the SacreBLEU implementation for both scores. Together, these metrics provide a strong comparison between the machine-translated text and human-translated reference. To normalize both scores to the same scale, we use the formula $score=\frac{100-HTER}{200}+\frac{chrF3}{100}$
@@ 5 Conclusions
GEPA optimization elevates minimal DSPy signatures, often closing the gap to expert-level performance, but the degree and direction of remaining differences are task-dependent. In terminology insertion, expert and optimized prompts are largely statistically indistinguishable on translation quality (BLEU, HTER), with optimization producing significantly higher term match rates for several models but not consistently improving fluency. In translation, most model–prompt differences are not statistically significant, with the exception of Gemini 3.1 Flash-Lite, where expert prompts significantly outperform optimized ones on BLEU. In LQA, optimization yields statistically significant gains over manual prompts on error characterization for several models, more consistently for Gemini 3.1 Flash-Lite. The degree to which optimization improves over base DSPy signatures also varies by task: gains are largest and most consistent in translation and LQA, while terminology insertion shows smaller improvements. The magnitude of optimization gains also varies by model: Qwen3:8B-q4_K_M, while performing worse than proprietary models on translation and terminology insertion, narrowed the gap on several LQA metrics, with optimization over base DSPy producing the largest single gain observed across all tasks and models, suggesting optimization may yield larger returns over base DSPy when the model is weaker.
@@
We highlight three concrete takeaways for practitioners. First, unified single-stage systems generally match or outperform decomposed multi-stage pipelines—most clearly for error detection in LQA (0.05–0.20 points on Detection F1) and edit distance in terminology insertion—suggesting end-to-end designs should be preferred unless there is a specific reason to decompose. Second, GEPA reliably elevates minimal DSPy signatures across tasks and models, making it a strong default when labeled data is available. Third, within unified systems, expert prompts retain an advantage for LQA error detection while optimization helps characterization —no single approach wins on every sub-metric. This suggests that combining the two, for example by seeding GEPA with expert prompts and letting it refine instructions, may be especially valuable for LQA. Practitioners should choose between approaches, or combine them, based on their available resources (labeled data, human effort, compute) and task requirements, rather than treating one as inherently superior.
@@
This study has several limitations. First, the comparison involves an inherent asymmetry in how each approach uses labeled data. GEPA optimization programmatically exploits gold-standard training and validation splits through automated feedback loops, whereas expert prompt engineering draws on domain knowledge, iterative testing, and, where available, selective inspection of examples. Second, the two approaches are not mutually exclusive: a domain expert could use GEPA optimization as a starting point and manually refine the resulting prompts, or conversely, use expert-crafted prompts as seed inputs for automatic optimization. Our study treats them as independent conditions, but a combined workflow may outperform either in isolation. Third, several design choices were scoped for practical reasons: we evaluate a single optimizer (GEPA) in its “light” mode with the dspy.Predict module, we do not evaluate few-shot or chain-of-thought manual prompts, and all evaluation is automatic without human judgments. Each of these choices may underestimate the ceiling of either approach: stronger optimizers, richer expert baselines, and human evaluation could all shift the results, but they were necessary to keep the study tractable across three tasks, five models, and multiple conditions. Fourth, all language pairs are English-centric, reflecting the availability of gold-standard data per task; results may not generalize to non-English-centric directions.